How do I solve this equation using the subtitution method:


60+0.20(x-550)=120+0.15(x-800)
1
Expert's answer
2011-07-12T05:32:57-0400
60+0.20(x-550)=120+0.15(x-800)
120 - 60 = 0.2x - 110 - 0.15x + 120
50 =0.05x
x = 1000